OpenStudy (jazilove):

I NEED HELP FAST- WILL MEDAL Given triangle ABC with A(-4, -2), B(4, 4), and C(18, -8), write the equation of the line containing the median that passes through point C in slope-intercept form.

OpenStudy (anonymous):

first, you have to find the midpoint of the opposing side: \[x = \frac{ -4 + 4 }{ 2 } = 0\] \[y = \frac{ -2 + 4}{ 2 } = 1\] so now we have the midpoint of segment AB, the opposing side. next we find the equation of the line that passes through C and this midpoint: \[(18, -8)(0, 1)\] \[m = \frac{ 1 - (-8)}{ 0 - 18 } = \frac{ 9 }{ -18 } = \frac{ -1 }{ 2 }\] now we have the slope of the median, and since the midpoint (0, 1) happens to also be the intercept, we can state the median in slope-intercept form: \[y = \frac{ -1 }{ 2 }x + 1\] and there you have it.

OpenStudy (jazilove):

you are seriously awesome, thank you!

OpenStudy (anonymous):

You're always welcome